You want to retire with $370 000 in the bank and you are able to earn 4.84% compounded quarterly for the next 25 years. You currently have $175 000 saved up. How much more do you have to invest in 15 years in order to achieve your goal?

A) $100 649.58

B) $110 649.58

C) $120 649.58

D) $130 649.58

E) $0.00
Answer #1

Rate per quarter = 4.84%/4 = 1.21%
Number of Periods = 4*25 = 100
The FV of 175000 = 175000*(1+1.21%)^100 = 582,613.4877

Since FV of saved amount is greater than 370,000 hence amount required is 0.

Option E.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well